OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = 2. P(x) = 2x^3 - 9x^2 + 27 P(2) = _____ Enter only a number.

OpenStudy (anonymous):

nevermind , i got it

OpenStudy (campbell_st):

just substitute x = 2 into the polynomial P(2) = 2*(2)^3 - 9*(2)^2 + 27 P(2) = 16 - 36 + 27 P(2) = 7